Circular No. 7664

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001de IN UGC 12089 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong, Rolvenden, Kent, of an apparent supernova (mag 18.0) on unfiltered CCD images secured for the U.K. Nova/Supernova Patrol on July 13.967 UT with a 0.30-m Schmidt-Cassegrain telescope. The following precise position by Armstrong is the average of two images: α = 22^{h}34^{m}18^{s}.06, δ = +50^{o}24′54′′.7 (equinox 2000.0), which is 1′′ west and 10′′ north of the center of UGC 12089. Confirming images were obtained by R. Arbour, South Wonston, U.K., on July 15.935 and by Armstrong on July 15.949 (when SN 2001de had showed brightening to mag 17.4). Armstrong adds that the new object is absent from his own CCD image taken on 2000 Sept. 23 (limiting mag 19.5), and it is absent from Palomar Sky Survey images taken on 1987 Oct. 17 (blue, limiting mag 22.5), 1989 Aug. 27 (red, limiting mag 20.8), and 1990 Oct. 28 (infrared).

SUPERNOVA 2001da IN NGC 7780 M. Dennefeld, Institut d’Astrophysique de Paris; A. Pasquali, Euro-

pean Southern Observatory, Garching; and J. M. Desert, F. McGroarty, S. Marchi, and J. Nirski, School of the Network of European Observatories in the North, Haute-Provence, report spectroscopy (range 370–730 nm, resolution 0.6 nm) of SN 2001da (cf. IAUC 7658) at Haute-Provence with the 1.93-m telescope (+ Carelec spectrograph) on July 13.10 UT. The shape of the continuum, strength of the blue Ca II/Si II features, a marked Si II 635-nm absorption, and no detectable H or He features indicate a type-Ia supernova close to maximum. With a recession velocity of the parent galaxy of 5155 km/s (Lyon Extragalactic Database), the red Si II absorption gives an expansion velocity of 12 000 km/s.

R. Chornock, M. Modjaz, and A. V. Filippenko, University of California at Berkeley, report that inspection of a CCD spectrum (range 330–1000 nm), obtained on July 16 UT with the Shane 3-m reflector at the Lick Observatory, reveals that SN 2001da is a type-Ia supernova. The spectralfeature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 0 ± 2 days relative to maximum light.

SUPERNOVA 2001dd IN UGC 11579 W. M. Wood-Vasey, Lawrence Berkeley Laboratory, reports unfiltered

magnitudes of SN 2001dd, from observations with the NEAT/MSSS 1.2-m reflector at Haleakala: June 16.44 UT, [17.7; 30.57, 16.8; July 13.60, 16.9.

2001 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green